OpenStudy (disco619):

How do I simplify HUGE Fractions without Calculator? Please Explain how with detail!!! 1) 800000/3200000 2) 7846837/33628738

OpenStudy (phi):

you can "cancel" equal number of trailing zeros from the top and bottom, for starters

OpenStudy (anonymous):

take off the zeroes for the first one so u would get 8/32 simplify that even farther and u get 1/4

OpenStudy (phi):

for the bottom one, I would round to 8,000,000/34,000,000 cancel zeros 8/34 divide by 2 4/17 or even more roughly, 4/16 = 1/4

OpenStudy (anonymous):

do they want it exactly?

OpenStudy (disco619):

So we have to remove the zeroes till they are equal in both numbers? Example: 100[00000000] & 1[00000000] which will make: 100 & 1 ????????

OpenStudy (anonymous):

which i doubt but u never know i suppose

OpenStudy (anonymous):

and yea that works take off zeros of equal amounts from the top and bottom and from there simplify it

OpenStudy (disco619):

Is THIS: 7846837/33628738 Possible without calculator?

OpenStudy (phi):

yes. 100/10 is the same as 10 or 1000/100 is 10 you are dividing top and bottom by 10 each time you cancel a 0

OpenStudy (anonymous):

depends do u have to round or get an exact answer?

OpenStudy (disco619):

Actually the real question was: 800000/3200000 But i wrote this: 7846837/33628738 myself to see how to solve it without that "trailing zero" process

OpenStudy (anonymous):

why.....

OpenStudy (anonymous):

eh the 2nd one would take forever to solve if you had to solve as is

OpenStudy (disco619):

Cause I have an Exam today and Im "sheit" in Maths. Don't even know how to simplify Fractions... And the Exam is without Calculator

OpenStudy (disco619):

Oh OK then, that means they would not give that kinda Question in exam which has Calculator Prohibited right?

OpenStudy (anonymous):

just find a number that they can both b divided by in most cases 2 is easier but takes longer

OpenStudy (phi):

notice that 7846837/33628738 =0.23333724268 and the rounded estimate 4/17 = 0.235

OpenStudy (anonymous):

pretty much what he said

OpenStudy (disco619):

WITHOUT CALCULATOR IS THE SECOND ONE POSSIBLE? I MEAN WOULD THEY GIVE IN EXAM WHICH DOESN'T HAVE CALCULATOR INVOLVED?

OpenStudy (anonymous):

easier ways and yes if you know simple arithmatics

OpenStudy (disco619):

THE FIRST I UNDERSTAND

OpenStudy (phi):

if you have time, memorize some of the divisibility rules http://www.mathsisfun.com/divisibility-rules.html

OpenStudy (phi):

WITHOUT CALCULATOR IS THE SECOND ONE POSSIBLE? I MEAN WOULD THEY GIVE IN EXAM WHICH DOESN'T HAVE CALCULATOR INVOLVED? only if you are testing to be "rain man"

OpenStudy (disco619):

Wut?

OpenStudy (phi):

http://en.wikipedia.org/wiki/Rain_Man

OpenStudy (anonymous):

lol just do this if they actually do give you something like that take one of the numbers find a number you can divide into that number evenly and see if it goes the same with the others and just keep on doing this